OpenStudy (anonymous):

convert the ellipse to standard form: 9x^2+y^2+36x-4y+31=0

OpenStudy (anonymous):

I hate factoring...

OpenStudy (anonymous):

Haha. Start by gathering all x's and y's like before. I'm going to factor the x terms too.\[9(x^2+4x)+y^2-4y=-31\] Then complete the squares, remembering what every you add to the x's, you have to multiply by 9, then add that to the other side.\[9(x^2+4x+4)+(y^2-4y+4)=-31+4+(9 \times 4)\] Then factor the perfect squares...\[9(x+2)^2+(y-2)^2=9\] Then divide everything by 9 and you end up with your answer.

OpenStudy (anonymous):

so, (x+2)^2/1 + (y-2)^2/9 ? @rkfitz053

OpenStudy (anonymous):

=1. Nailed it. I gotta go though. Any other quick questions?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

Ok, good luck on the rest.